John is twice as old as his daughter. In four years he will be three times what his daughter's age was three years ago. How old are they both right now?

I think I have the correct idea of how to set up my equations with my variables, but the three years ago is messing me up. Thank you.

Answer by htmentor(1343)

Let J = John's age now
Let D = his daughter's age now
So right now J = 2D
In 4 years John will be J+4
His daughter's age 3 years ago is D-3
The equation is
J+4 = 3(D-3) [John 4 years from now is 3 times his daughter 3 years ago]
But since J = 2D, we have
2D + 4 = 3(D - 3)
Solve for D:
2D + 4 = 3D - 9
This gives D = 13
So right now, daughter is 13 and John is 26.